OpenStudy (anonymous):

find all solutions for the fallowing equation sinxcosx-2sinx=0

OpenStudy (anonymous):

Factor :) \[\huge [\sin(x)][\cos(x)-2]=0\]

OpenStudy (anonymous):

okay you factored out the sin. now do i set the sinx and cosx-2 equal to zero?

OpenStudy (anonymous):

Yes. But cos(x) - 2 = 0 cos(x) = 2 ... cosine can never be equal to 2, so scratch that, your only chance is for sin(x) to be 0 :)

OpenStudy (anonymous):

okay i understand that. but when sin equals zero does that mean the answer is \[2piK\]

OpenStudy (anonymous):

where k is an integer?

OpenStudy (anonymous):

Actually, just \[\large k\pi \ \ \ \ \ \ k\in \mathbb{Z}\]

OpenStudy (anonymous):

Remember that \(\sin \pi = 0\) too.